kekeman:

Given ΔABC with ∠ B= 85°, ∠C =47°, and a = 19 inches, what is the length of b? A.) 11.577 inches B.) 14.321 inches C.) 20.108 inches D.) 25.470 inches

Vocaloid:

sketch of the triangle. not to scale. I've taken the liberty of calculating angle A by subtracting 180 - angle B - angle C.  set up the law of sines, sin(A)/a = sin(B)/b --> sin(48)/19 = sin(85)/b solve for b

kekeman:

@vocaloid wrote:
sketch of the triangle. not to scale. I've taken the liberty of calculating angle A by subtracting 180 - angle B - angle C. Created with RaphaëlABC8547 dd48 da19bcReply Using Drawing set up the law of sines, sin(A)/a = sin(B)/b --> sin(48)/19 = sin(85)/b solve for b
Length of b = 25.470

Vocaloid:

good